# Semi-orthogonality in Fukaya-Seidel mirrors to blowups of abelian varieties

### v1

@@ 3.2.4. Inverses: continuation maps
Since we will see in Claim 3.18 that the input and output in morphisms wrap in opposite directions, we need an inverse to quasi-units to connect Floer complexes in Lemma 3.14. These will be the continuation maps. The algorithm to define and use them is as follows.
@@
    Define the isotopies in the moving Lagrangian boundary condition. We use those induced by Equation (3.3).
@@
      Absorb the moving Lagrangian boundary condition into a perturbation to the Cauchy-Riemann equation [Aur14, Remark 1.10], known as Floer’s equation. Here it is
      (3.17) $(du+(\xi^{\tau})^{\#}\otimes d\tau)^{0,1}$
      for a suitable $\tau$ depending on two parameters. Continuation maps count these solutions.
@@
      Rescale the curve to obtain a solution to the Cauchy-Riemann equation.
  2. (2)
    Deduce results for elliptic operators which are needed to show that the continuation map moduli space in Equation (3.23) can be counted.
    1. (a)
      A 0-dimensional manifold: A regular $J$ exists, meaning the linearization of the Cauchy-Riemann operator at each $u$ in the moduli space is surjective. This is a standard argument provided curves are shown to have a dense set of somewhere injective points.
    2. (b)
      Compact: this follows from Gromov compactness. This is a standard argument provided there’s a maximum principle that the curves in the moduli space stay in a compact subset for a given homology class $[u]$.
We follow the methodology of [AA24, Chapter 3], noting that we have no $X_{H}$ term since we have no wrapping in the fiber. Another related setting is that of [GPS24, §8.6], since curves projected to the base are then in an exact symplectic manifold.
@@
This isotopy moves in the opposite direction of the moving boundary condition on the quasi-unit. Because $L^{j}=\rho^{-\tilde{\epsilon}j}(L)$, the isotopy $\Lambda_{s}$ moves $L^{j+1}$ to $L^{j}$. That is, there exists $\tilde{K}$ so that
(3.19) $\Lambda_{s}(L^{j})=\begin{cases}L^{j+1},&s<-\tilde{K}\\ L^{j},&s>\tilde{K}\end{cases}$.
So in words, $\Lambda_{s}(L):\mathbb{R}\to Y^{0}$ rotates Lagrangian ends of $W_{0}(L)$ counterclockwise outside a compact set around 0, meaning $\frac{d}{ds}\theta_{h}(\Lambda_{s}(L))>0$ for $\theta_{h}$ as in Definitions 3.4 and 3.6. When $L=K_{j}$, $W_{0}(\Lambda_{s}(L))$ also fully wraps clockwise around 0. We cancel this movement by introducing a perturbation term moving in the direction of $\rho^{\tilde{\epsilon}\tilde{\chi}(s)}$ on the boundary:
@@
Thus $\tau$ is a way to extend the Lagrangian perturbation on the boundary of the disc to the whole disc, providing enough parameters in the perturbation to ensure transversality, while also canceling the moving Lagrangian boundary condition in $\Lambda_{s}(L)$. It is valued in the vector $(\xi^{\tau})^{\#}$ on the target and $d\tau$ inputs vectors on the domain, so tensored together it is the same type as $du$. It describes how $du$ should be perturbed infinitesimally over the domain of $u$.
To ensure enough perturbation parameters to prevent solutions to the Floer equation that have a lot of symmetry because they are nowhere somewhere injective, we need the full set of two parameters on the domain $\Sigma$ in both $s$ and $t$. We also perturb with the correct orientation for a maximum principle to hold, that is, in directions out of the strip on the boundary. This is related to [Sei08, §(8k), Equation (8.21)] for moving boundary conditions as in continuation maps; in cases such as for a differential on $CF(K_{j},K_{j})$, one can perturb in the $t$-direction only. That is, one can use $\xi^{\#}\otimes dt$ near 0 so the perturbation is 0 on $\partial_{s}$ tangent to the boundary and $\xi^{\#}$ on $\partial_{t}$ normal to the boundary.
@@
(3.22) $F_{L_{0}^{k},L_{1}^{j}}(p)=\sum_{q\in L_{0}^{k+1}\cap L_{1}^{j+1}}\sum_{\begin{subarray}{c}[u]:[u]=\beta\in H_{2}(Y^{0},\cup_{s\in\mathbb{R}}\Lambda_{s}(L_{0}^{k})\bigcup\cup_{s\in\mathbb{R}}\Lambda_{s}(L_{1}^{j}))\\ \text{ind}(u)=0\end{subarray}}\#\mathcal{M}(p,q;[u],J)e^{-2\pi\int_{\beta}\omega}\cdot q$
(3.23) $\mathcal{M}(p,q;[u],J)$ $=\{u:(\mathbb{R}\times[0,1],\mathbb{R}\times\{0\}\cup\mathbb{R}\times\{1\})\to(Y^{0},\cup_{s\in\mathbb{R}}\Lambda_{s}(L_{0}^{k})\bigcup\cup_{s\in\mathbb{R}}\Lambda_{s}(L_{1}^{j})\mid$
@@
Theorem 3.10.
The moduli space $\mathcal{M}(p,q;[u],J)$ is a (1) smooth manifold, (2) compact, and (3) 0-dimensional, so it is a finite set of points which can be counted.
Proof.
These arguments are standard in the literature ([AD14, §8] for moduli spaces of cylinders between Reeb orbits, [MS12, §6.7] for domain dependent $J$, and [Sei08, §(9k), Equation (9.27)], [AS10, Lemma 8.5], [Sei12, §4] for discs). We adapt them to our setting. For 3), the moduli space is 0-dimensional because we count index 0 discs.
(1) The moduli space is the vanishing of the differential operator $u\mapsto(du+(\xi^{\tau})^{\#}\otimes d\tau)^{0,1}$, thought of as a section of a Banach bundle, so by the Sard-Smale theorem it suffices to show the derivative (also known as the linearization) of the differential operator at each map $u$ in the moduli space, is surjective. (A finite-dimensional example of this process is the vanishing of $f(x,y,z)=x^{2}+y^{2}+z^{2}-a:\mathbb{R}^{3}\to\mathbb{R}$, which cuts out a smooth manifold for $a\neq 0$, a sphere, because its linearization $df=[2x\;2y\;2z]\neq\underline{0}$ at each such $(x,y,z)\in f^{-1}(0)$.)
We have two choices of datum, one for $\tau$ and one for $J$. The linearization of $(du+(\xi^{\tau})^{\#}\otimes d\tau)^{0,1}$ is
@@
where the first two terms are justified in [Can20, Equation (4.15)] and the third term is from an infinitesimal perturbation of $\tau$.
Remark 3.11.
In the notation of [AD14, §8.4], their perturbation is $JX(s,t)$ so they linearize
$\mathcal{F}(u+Y)=\frac{\partial(u+Y)}{\partial s}+J(u+Y)\frac{\partial(u+Y)}{\partial t}-J(u+Y)X_{t}(u+Y)$
to the linear components
(3.24) $\frac{\partial Y}{\partial s}+(dJ)_{u}(Y)\frac{\partial u}{\partial t}+J(u)\frac{\partial Y}{\partial t}-(dJ)_{u}(Y)X_{t}-J(u)(dX_{t})_{u}(Y).$
In this paper, the perturbation term comes from $\tau(s,t)$ so instead of the last two terms in Equation (3.24) we have a single $\dot{\tau}^{0,1}$ term, what Seidel refers to as $(\delta Y)^{0,1}$ in [Sei08, Equation (9.26)], see also [AS10, Equation (8.10)] for perturbation term $X\otimes\gamma$ and linearization with respect to $u$ and $J$.
@@
Recall that an injective point $z\in\mathbb{D}$ is one such that $d_{z}u\neq 0$ and $u^{-1}(u(z))=\{z\}$. We make use of the perturbation to ensure that discs are somewhere injective. Decomposing discs with Lagrangian boundary condition to somewhere injective discs has been studied in [KO00, Laz00, Laz11]. Here we show discs are already somewhere injective.
@@
Discs aren’t multiply covered because $\tilde{u}:=W_{0}\circ u:\mathbb{R}\times[0,1]\to\mathbb{C}$ is an unramified holomorphic map of Riemann surfaces. The two boundary components map to different moving Lagrangians which prevents ramification points on the interior. A ramification point on the boundary means $d\tilde{u}$ collapses $\partial_{s}$ and $\partial_{t}=j\partial_{s}$ together but pseudo-holomorphicity implies one maps to the tangent space $TW_{0}(L)$ and the other to $J(TW_{0}(L))$. For a totally real submanifold, the only vector in their intersection $TW_{0}(L)\cap J(TW_{0}(L))=\{0\}$. Thus there couldn’t have been a ramification point on the boundary. So discs are not multiply covered, and now we justify why the set of injective points is dense.
To show non-injective points are isolated, first select $u:\Sigma\to Y^{0}$ a curve in the moduli space in Equation (3.23). Suppose by contradiction that we can pick a $z_{0}\in\Sigma$ such that there exists $z_{0}\neq z_{0}^{\prime}$ in $\Sigma$ and disjoint open neighborhoods of each, call them $U,U^{\prime}$ so that there is a biholomorphism $\varphi:U\to U^{\prime}$ where $u(z)=u\circ\phi(z)$ for all $z\in U$ but $z\neq\varphi(z)\in U^{\prime}$. Since $u|_{U}=u\circ\phi$ (and so $du|_{U}=du\circ d\phi$) agree on the neighborhoods $U,U^{\prime}$, but there exist vector $V$ and point $z$ where $(\xi^{\tau(z)})^{\#}d\tau_{z}(V)\neq(\xi^{\tau(\varphi(z))})^{\#}d\tau_{\varphi(z)}(d\varphi(V))$, then $u$ couldn’t have solved the Floer equation in Equation (3.23). The reason we introduced $\tau$ is to pick up the fact that $z\neq\phi(z)$ differ in the domain, even though from the perspective of the image they are the same. Note that $\tau$ cannot be constant by the definition of $\tilde{\chi}$ in Definition 3.8 and the definition of $\tau$ on the boundary by Equation (3.21). Thus non-injective points are isolated and injective points are dense. ∎
2) It remains to show the moduli space is compact. Since the moduli space of discs is a metric space, it suffices to show it is sequentially compact. Thus we need to show that every sequence has a convergent subsequence whose limit is in the space. This is done using Gromov compactness. To satisfy the assumptions of, and then apply, Gromov compactness, we split into two parts.
@@
    We need to exclude discs escaping to infinity. Then Gromov compactness applies and states there are only certain types of bubbling that could be in the limit.
@@
    We exclude those possible degenerations of curves.
2a) To exclude discs escaping to infinity, we need a maximum principle. Consider the annulus $A$ from Equation (3.4). Outside of a larger annulus, projected Lagrangian ends wrap according to the Hamiltonian in Equation (3.3).
@@
For $u\in\mathcal{M}(p,q;[u],J)$, the image of the disc $u$ is contained in a compact subset of $Y^{0}$.
@@
First, let’s define the region we expect the curve to stay in within the base, because fibers are already compact. Enlarge $A$ to a larger annulus $\tilde{A}:=\{r_{1}\leq|w|\leq r_{2}\}\ni W_{0}(p),W_{0}(q)$ containing the points the projected strip limits to, and also $(R_{1},R_{4})\subset(r_{1},r_{2})$ so outside of $\tilde{A}$ projected Lagrangians have constant angle and isotopies only rotate them by Equation (3.3). We’d like to show $u(\mathbb{R}\times[0,1])$ is contained in the compact subset $\Omega:=Y^{0}|_{W_{0}^{-1}(\tilde{A})}$. We will do this by showing that maximums of $|W_{0}\circ u|$ and $|\frac{1}{W_{0}}\circ u|$, if they exist, must be attained on the boundary. Then we show that they cannot be attained on the boundary $\mathbb{R}\times\{0\}\cup\mathbb{R}\times\{1\}$ outside $\Omega$ and $1/\Omega$ respectively. Therefore they must be attained in $\Omega$ or a maximum is approached at the limits to $p$ and $q$, both of which are contained in $\Omega$.
Recall that away from the singular fiber $J|_{Y^{0}\backslash\Omega}=J_{0}$ is multiplication by $i$ in the affine toric coordinates $t_{1},t_{2},t_{3}$ which are related to $y_{1},y_{2},y_{3}$ of Remark 1.4 by the process described in [ACL21, Equation (72)]. In particular, since $v_{0}=t_{3}$ we have $W_{0}=T^{\epsilon}(t_{3}-1)$ and so $W_{0}|_{\Omega}$ is $(J_{0},i)$-holomorphic because $dW_{0}\circ J_{0}=T^{\epsilon}dt_{3}\circ i=i\circ d(T^{\epsilon}(t_{3}-1))=i\circ dW_{0}$. Similarly $\frac{1}{W_{0}}:\mathbb{R}\times[0,1]\to\mathbb{C}^{*}$ is $(J_{0},i)$ holomorphic by composing with the $i$-holomorphic map $w\mapsto 1/w$.
@@
Let $j_{1}:=j$ and $j_{2}:=(\rho^{\tau})_{*}j$ so $\tilde{u}$ is $(j_{1},j_{2})$-holomorphic on domains in $\mathbb{C}$, thus in particular it is conformal for the metrics $g_{1},g_{2}$ compatible with $\omega_{\mathbb{C}}$ defined by $g_{k}(-,-)=\omega_{\mathbb{C}}(-,j_{k}(-))$ for $k=1,2$. Let $\tilde{u}:=r(s,t)e^{i\theta(s,t)}$ so $r(s,t)=r\circ\tilde{u}$ and $\theta(s,t)=\theta\circ\tilde{u}$. Then by expressing $d\tilde{u}$ in real coordinates $\tilde{u}=:\tilde{u}_{x}+i\tilde{u}_{y}$ and converting to polar coordinates, using Equation (3.25) we have
(3.26) $d\tilde{u}(\partial_{s})=-j_{2}\circ d\tilde{u}\circ j_{1}(\partial_{s})\implies\partial_{s}r(s,t)+ir(s,t)\partial_{s}\theta(s,t)=-j_{2}\partial_{t}r(s,t)-j_{2}ir(s,t)\partial_{t}\theta(s,t).$
We need to check that the disc boundary doesn’t stretch out infinitely along the Lagrangians under $\tilde{u}$. We start by showing boundedness away from infinity. Suppose there’s a maximum $r(s_{0},t_{0})$ on the boundary and outside of $\tilde{A}$, at some $(s_{0},t_{0})\in W_{0}(Y^{0})\backslash\{|w|\leq r_{2}\}$ where $t_{0}$ is 0 or 1. Then $\partial_{s_{0}}(r\circ\tilde{u})=0$ there. By the Hopf lemma, at a maximum, the derivative in the other direction $\partial_{t}$ moving outside the disc, must be positive. Because $e^{\pi z}$ is a biholomorphism from $\mathbb{R}\times(0,1)$ to the interior of the upper-half plane which extends to mapping $\mathbb{R}\times\{0\}\to\mathbb{R}_{+}$ and $\mathbb{R}\times\{1\}\to\mathbb{R}_{-}$, we can take coordinates on the strip to be parametrized by points on the upper-half plane (UHP) and then $\partial_{t^{UHP}}$ points outwards of the upper-half plane where $s^{UHP}+it^{UHP}$ are the UHP coordinates. The derivative of $r$ on $\tilde{u}$ must point outwards at $s_{0}^{UHP}+it_{0}^{UHP}:=e^{\pi(s_{0}+it_{0})}$ in the direction of $\partial_{t^{UHP}}$ so
$\frac{\partial(r\circ\tilde{u})}{\partial t^{UHP}}\mid_{(s_{0},t_{0})}<0.$

### v2

@@ 3.2.4. Inverses: continuation maps
Since we will see in Claim 3.21 that the input and output in morphisms wrap in opposite directions, we need an inverse to quasi-units to connect Floer complexes in Lemma 3.16. These will be the continuation maps. The algorithm to define and use them is as follows. We follow the methodology of [AA24, Chapter 3], noting that we have no $X_{H}$ term since we have no wrapping in the fiber. Another related setting is that of [GPS24, §8.6], since curves projected to the base are then in an exact symplectic manifold.
@@
    Define the isotopy of the moving Lagrangian boundary according to the wrapping needed. We use that in Equation (3.11) as a lift of the Hamiltonian vector field associated to the Hamiltonian in the base in Equation (3.9), since we have no wrapping in the fiber. The moving Lagrangian boundary condition introduces a perturbation into the Cauchy-Riemann equation [fuk_intro, Remark 1.10], known as Floer’s equation. Here, it is
    (3.24) $(du+(\xi^{\tau})^{\#}\otimes d\tau)^{0,1}=0$
    for a suitable $\tau$ depending on two parameters (different from the $\tau$ used to denote the complex parameter on $X$). Continuation maps count these solutions.
  2. (2)
    Show that the continuation map moduli space in Equation (3.30) can be counted.
@@
      A 0-dimensional manifold: By Fredholm theory, it suffices to show that a regular $J$ exists. This means that the linearization of the Floer operator in Equation (3.24) is surjective. This is a standard argument provided curves $u$ in the moduli space are shown to have a dense set of somewhere injective points, which follows from the choice of perturbation $\tau$.
@@
      Compact: this follows from Gromov compactness. This is a standard argument provided there’s a maximum principle that the curves $u$ in the moduli space stay in a compact subset for a given relative homotopy class $[u]$. This follows from rescaling the curve to obtain a solution to the elliptic Cauchy-Riemann equation $(du)^{0,1}=\overline{\partial}_{J}(u)=\frac{1}{2}(du+J\circ du\circ j)=0$ which satisfies a maximum principle on the interior, and then a boundary analysis for the boundary.
@@
This isotopy moves in the opposite direction of the moving boundary condition on the quasi-unit. Because $L^{k}=\rho^{-\tilde{\epsilon}k}(L)$, the isotopy $\Lambda_{s}$ moves $L^{k+1}$ to $L^{k}$. That is, there exists $\tilde{K}$ so that
(3.26) $\Lambda_{s}(L^{k})=\begin{cases}L^{k+1},&s<-\tilde{K}\\ L^{k},&s>\tilde{K}\end{cases}$.
So in words, $\Lambda_{s}(L):\mathbb{R}\to Y^{0}$ rotates Lagrangian ends of $W_{0}(L)$ counterclockwise outside a compact set around 0, meaning $\frac{d}{ds}\theta_{h}(\Lambda_{s}(L))>0$ for $\theta_{h}$ as in Definitions 3.3 and 3.5, by Equation (3.13). When $L=K_{j}$, $W_{0}(\Lambda_{s}(L))$ also fully wraps clockwise around 0 by Equation (3.16).
We introduce a perturbation term moving in the direction of $\rho^{\tilde{\epsilon}\tilde{\chi}(s)}$ on the boundary:
@@
Thus $\tau$ is a way to extend the Lagrangian perturbation on the boundary of the disc to the whole disc, providing enough parameters in the perturbation to ensure transversality of the moduli space of solutions to the perturbed equation in Equation (3.24), while also allowing us to cancel the moving Lagrangian boundary condition in $\Lambda_{s}(L)$ by rescaling the map $u$. The perturbation term $(\xi^{\tau})^{\#}\otimes d\tau$ is valued in the vector $(\xi^{\tau})^{\#}$ on the target and $d\tau$ inputs vectors on the domain, so tensored together it is the same type as $du$. It describes how $du$ should be perturbed infinitesimally over the domain of $u$.
To ensure enough perturbation parameters to prevent solutions to the Floer equation that have a lot of symmetry because they are nowhere somewhere injective, we need $\tau$ to depend on the full set of two parameters on the domain $\Sigma$, so both $s$ and $t$. We also perturb with the correct orientation for a maximum principle to hold, that is, in directions out of the strip on the boundary. This is related to [seidel, §(8k), Equation (8.21)] for moving boundary conditions as in continuation maps. (In cases such as for a differential on $CF(W_{0}(K_{j}),W_{0}(K_{j}))$, one can perturb in the $t$-direction only. That is, one can use $(\xi^{\tau})^{\#}\otimes dt$, which is 0 on $\partial_{s}$ tangent to the boundary and $(\xi^{\tau})^{\#}$ on $\partial_{t}$ normal to the boundary.)
@@
(3.29) $F_{L_{0}^{k},L_{1}^{j}}(p)=\sum_{q\in L_{0}^{k+1}\cap L_{1}^{j+1}}\sum_{\begin{subarray}{c}[u]:[u]=\beta\in\pi_{2}(Y^{0},\cup_{s\in\mathbb{R}}\Lambda_{s}(L_{0}^{k})\bigcup\cup_{s\in\mathbb{R}}\Lambda_{s}(L_{1}^{j}))\\ \text{ind}(u)=0\end{subarray}}\#\mathcal{M}(p,q;[u],J)e^{-2\pi\int_{\beta}\omega}\cdot q$
(3.30) $\mathcal{M}(p,q;[u],J)$ $=\{u:(\mathbb{R}\times[0,1],\mathbb{R}\times\{0\}\cup\mathbb{R}\times\{1\})\to(Y^{0},\cup_{s\in\mathbb{R}}\Lambda_{s}(L_{0}^{k})\bigcup\cup_{s\in\mathbb{R}}\Lambda_{s}(L_{1}^{j}))\mid$
@@
Lemma 3.10.
The moduli space $\mathcal{M}(p,q;[u],J)$ is a smooth 0-dimensional manifold.
Lemma 3.11.
The moduli space $\mathcal{M}(p,q;[u],J)$ is compact.
Corollary 3.12.
The moduli space $\mathcal{M}(p,q;[u],J)$ is a finite set of points which can be counted.
Proof of Corollary 3.12.
By Lemmas 3.10 and 3.11, the moduli space is a compact 0 dimensional manifold. ∎
Proof of Lemma 3.10.
This is a standard result in the literature ([AD14, §8] for moduli spaces of cylinders between Reeb orbits, [jholbk, §6.7] for domain dependent $J$, and [seidel, §(9k), Equation (9.27)], [AS10, Lemma 8.5], [Se12, §4] for discs). We apply them to our setting. The moduli space is 0-dimensional because we count index 0 discs.
The moduli space is the vanishing of the differential operator $u\mapsto(du+(\xi^{\tau})^{\#}\otimes d\tau)^{0,1}$, thought of as a section of a Banach bundle, so by the Sard-Smale theorem it suffices to show the derivative (also known as the linearization) of the differential operator at each map $u$ in the moduli space, is surjective. (A finite-dimensional example of this process is the vanishing of $f(x,y,z)=x^{2}+y^{2}+z^{2}-a:\mathbb{R}^{3}\to\mathbb{R}$, which cuts out a smooth manifold for $a\neq 0$, a sphere, because its linearization $df=[2x\;2y\;2z]\neq\underline{0}$ at each such $(x,y,z)\in f^{-1}(0)$.) We have three choices of datum to vary: $u$, $J$, and $\tau$. The linearization of $(du+(\xi^{\tau})^{\#}\otimes d\tau)^{0,1}$ is
@@
where the first two terms are justified in [Ca20, Equation (4.15)] and the third term is from an infinitesimal perturbation of $\tau$. (In [AD14, §8.4] their Floer equation involves a perturbation from $JX(s,t)$ instead of $\tau$. Our $\dot{\tau}^{0,1}$ is similar to $(\delta Y)^{0,1}$ in [seidel, Equation (9.26)], and see [AS10, Equation (8.10)] for perturbation term $X\otimes\gamma$ and linearization with respect to $u$ and $J$.)
@@
Recall that an injective point $z\in\mathbb{D}$ is one such that $d_{z}u\neq 0$ and $u^{-1}(u(z))=\{z\}$. We make use of the perturbation $\tau$ to ensure that discs are somewhere injective. Decomposing discs with Lagrangian boundary condition to somewhere injective discs has been studied in [KO00, lazz1, lazz2]. Here we show discs are already somewhere injective. Note that the boundary conditions prevent multiply covered discs.
@@
To show non-injective points are isolated, first select $u:\Sigma\to Y^{0}$ a curve in the moduli space in Equation (3.30). Suppose by contradiction that we can pick a $z_{0}\in\Sigma$ such that there exists $z_{0}\neq z_{0}^{\prime}$ in $\Sigma$ and disjoint open neighborhoods of each, call them $U,U^{\prime}$ so that there is a biholomorphism $\varphi:U\to U^{\prime}$ where $u(z)=u\circ\varphi(z)$ for all $z\in U$ but $z\neq\varphi(z)\in U^{\prime}$. Because $u|_{U}=u\circ\varphi$, we have $du|_{U}=du\circ d\varphi$ on $U$. But there exist vector $v$ and point $z\in U$ where $(\xi^{\tau(z)})^{\#}d\tau_{z}(v)\neq(\xi^{\tau(\varphi(z))})^{\#}d\tau_{\varphi(z)}(d\varphi(v))$. Then $u$ couldn’t have solved the Floer equation in Equation (3.24) at $\varphi(z)$ if it did at $z$. The perturbation $\tau$ distinguishes that $z\neq\varphi(z)$ in the domain, even though from the perspective of the image of $u$ they are the same. Thus non-injective points are isolated and injective points are dense. ∎
This concludes the proof of Lemma 3.10 that $\mathcal{M}(p,q;[u],J)$ is a smooth 0-dimensional manifold. ∎
Now we prove Lemma 3.11 that the moduli space is compact. Since the moduli space of discs is a metric space [jholbk, p47], it suffices to show it is sequentially compact, that every sequence has a convergent subsequence whose limit is in the space. This is ensured by Gromov compactness, named after [Gr85]. To prove Lemma 3.11, we split into two parts.
@@
    We exclude discs escaping to infinity, in Lemma 3.15, by proving a maximum principle. Then Gromov compactness applies and states there are only certain types of bubbling that could be in the limit, Proposition 3.14.
@@
    We exclude the possible degenerations of curves, in the proof of Lemma 3.11.
Proposition 3.14.
A sequence of pseudo-holomorphic discs with Lagrangian boundary conditions, and satisfying a maximum principle, has a convergent subsequence, which limits to a pseudo-holomorphic curve that may degenerate. Possible degenerations are
  1. (1)
    sphere bubbles (energy concentrates at an interior point)
  2. (2)
    disc bubbles (energy concentrates at a boundary point)
  3. (3)
    strip-breaking (energy concentrates at a limiting intersection point).
Proof.
This is proved for compact symplectic manifolds and discs with Lagrangian boundary conditions in [jholbk, §4.6, Lemma 4.6.5], which applies once we have the maximum principle in a compact set. See also [fooo1, Example 2.4.14, Theorem 2.4.36]. ∎
Consider the annulus $A$ from Equation (3.10). Outside of $A$, projected Lagrangian ends wrap positively under $\rho^{t}$ for increasing $t$. Recall that in an LG model, curves counted in structure maps are sections of the superpotential $W_{0}$ [seidel, Equation (17.2)]. In other words, because $u$ is 1-1 there exists $z(s,t)\in\mathbb{D}$ uniquely defined by
(3.31) $(W_{0}\circ u)(z(s,t))=A(s,t)$
where $A(s,t)$ parametrizes
(3.32) $D:=(W_{0}\circ u)(\mathbb{D}).$
Then we can define $u_{D}:D\to Y^{0}$
(3.33) $u_{D}(A(s,t))=u(z(s,t))$
as a $((W_{0}\circ u)_{*}j,J)$-holomorphic section of $W_{0}$ if $u$ was a $(j,J)$-holomorphic curve.
@@
For all $u\in\mathcal{M}(p,q;[u],J)$, the image of $u$ is contained in a compact subset of $Y^{0}$.
@@
By Equation (3.33), we may assume $u$ is a section over a bigon in the base of $W_{0}$, and we identify $(s,t)\in\mathbb{R}\times[0,1]$ with $A(s,t)\in\mathbb{C}^{*}$. First, let’s define the region we expect the curve to stay in within the base, because fibers are already compact. Enlarge $A$ to a larger annulus $\tilde{A}:=\{\tilde{R}_{1}\leq|w|\leq\tilde{R}_{2}\}\ni W_{0}(p),W_{0}(q)$ containing the points the projected strip limits to. Also let $(R_{1},R_{4})\subset(\tilde{R}_{1},\tilde{R}_{2})$ so outside of $\tilde{A}$, projected Lagrangians have constant angle and isotopies only rotate them by Equation (3.9). We’d like to show $u(\mathbb{R}\times[0,1])$ is contained in the compact subset $\Omega:={W_{0}^{-1}(\tilde{A})}$. We will do this by showing that maximums of $|W_{0}\circ u|$ and $|\frac{1}{W_{0}}\circ u|$, if they exist, must be attained on the boundary. Then we show that any boundary maximum, if it exists, must lie within $\Omega$. If it doesn’t exist, then a maximum is approached as we limit to $p$, $q$, and their images are also in $\Omega$. Therefore the image of $u$ is contained in $\Omega$.
Recall that away from the singular fiber, $J|_{Y^{0}\backslash\Omega}=J_{0}$ is multiplication by $i$ in the affine toric coordinates $t_{1},t_{2},t_{3}$ which are related to $y_{1},y_{2},y_{3}$ of Remark 1.3 by the process described in [ACL21, Equation (72)]. In particular, since $v_{0}=t_{3}$ we have $W_{0}=T^{\epsilon}(t_{3}-1)$ and so $W_{0}|_{\Omega}$ is $(J_{0},i)$-holomorphic because $dW_{0}\circ J_{0}=T^{\epsilon}dt_{3}\circ i=i\circ d(T^{\epsilon}(t_{3}-1))=i\circ dW_{0}$. Similarly $\frac{1}{W_{0}}:\mathbb{R}\times[0,1]\to\mathbb{C}^{*}$ is $(J_{0},i)$ holomorphic by composing with the $i$-holomorphic map $w\mapsto 1/w$.
@@
Let $j_{1}:=j$ and $j_{2}:=(\rho^{\tau})_{*}j$ so $\tilde{u}$ is $(j_{1},j_{2})$-holomorphic on domains in $\mathbb{C}$. In particular, it is conformal for the metrics $g_{1},g_{2}$ compatible with $\omega_{\mathbb{C}^{*}}$ defined by $g_{k}(-,-)=\omega_{\mathbb{C}^{*}}(-,j_{k}(-))$ for $k=1,2$. Let $\tilde{u}:=r(s,t)e^{i\theta(s,t)}$ so $r(s,t)=r\circ\tilde{u}$ and $\theta(s,t)=\theta\circ\tilde{u}$.
We need to check that the disc boundary doesn’t stretch out arbitrarily far along the Lagrangians under $\tilde{u}$. We start by showing boundedness away from infinity. Suppose there’s a maximum $r(s_{0},t_{0})$ on the boundary and outside of $\tilde{A}$, at some $(s_{0},t_{0})\in W_{0}(Y^{0})\backslash\{|w|\leq\tilde{R}_{2}\}$ where $t_{0}$ is 0 or 1. Then
(3.35) $\partial_{s_{0}}(r\circ\tilde{u})=0$
there. By the Hopf lemma, at a maximum, the derivative in the other direction $\partial_{t}$ moving outside the disc, must be positive. Because $e^{\pi(s+it)}$ is a biholomorphism from $\mathbb{R}\times(0,1)$ to the interior of the upper-half plane which extends to mapping $\mathbb{R}\times\{0\}\to\mathbb{R}_{+}$ and $\mathbb{R}\times\{1\}\to\mathbb{R}_{-}$, we can take coordinates on the strip to be parametrized by points on the upper-half plane (UHP) and then $\partial_{t^{UHP}}$ points outwards of the upper-half plane where $s^{UHP}+it^{UHP}$ are the UHP coordinates. The derivative of $r$ on $\tilde{u}$ must point outwards at $s_{0}^{UHP}+it_{0}^{UHP}:=e^{\pi(s_{0}+it_{0})}$ in the direction of $\partial_{t^{UHP}}$ so
(3.36) $\frac{\partial(r\circ\tilde{u})}{\partial t^{UHP}}\mid_{(s_{0},t_{0})}<0.$